OpenStudy (anonymous):

Solve by taking the a logarithm of both sides. 8^x = 23

OpenStudy (anonymous):

taking log on both sides we get xlog8=log23 x=log23/log8 x=1.50785

OpenStudy (anonymous):

no sumbul...you cant use log_10 without change base value! You could do what you did if you toke the log_8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

take ln then?

OpenStudy (anonymous):

we need to make log_c a log_b(a) = --------- log_c b

OpenStudy (anonymous):

all right

OpenStudy (anonymous):

@sumbul that's the same thing I did when I solved it.

OpenStudy (anonymous):

we may rewrite: 2^3x=23 taking log_2 on both sides: 3x= log_2 23 log_2 23 = (log_10 23) / (log_10 2) = 4,52 so 3x = 4.5235 x = 1.5078

OpenStudy (anonymous):

answers are same

OpenStudy (anonymous):

im was thinking about.. i was wrong when i said "no sumbul...you cant use log_10 without change base value! You could do what you did if you toke the log_8" We can use log_c a^b = b log_c a !! Sorry @sumbul =)